OpenStudy (anonymous):

please help log 0.07921+2log18.99-4log5.79-log.9474 =2(bar).8988+2.5572-3.0508-1(bar).9765 kindly help me in simplifying

OpenStudy (anonymous):

plz reply i m waiting

OpenStudy (mathstudent55):

log0.07921+ 2log18.99 - 4log5.79 - log.9474 = 2(bar).8988 + 2.5572 - 3.0508 - 1(bar).9765 = -2 + 0.8988 + 2.5572 - 3.0508 - (-1 + 0.9765) = -2 + 0.8988 + 2.5572 - 3.0508 + 1 - 0.9675

OpenStudy (anonymous):

thanks a lot

OpenStudy (mathstudent55):

wlcm